The sum of three numbers is 121. The third number is 4 times the first. The first number is 5 more than the second. What are the numbers ?
Found 2 solutions by Edwin McCravy, tajar:
Answer by Edwin McCravy(20059)

Let 
the second number = x

The first number is 5 more than the second.
The first number = x+5

The third number is 4 times the first.
The third number = 4(x+5)

The sum of three numbers is 121.
x + x+5 + 4(x+5) = 121

Solve and get x = 16

The first number = x+5 = 16+5 = 21
The second number = x = 16
The third number = 4(x+5) = 4(16+5) = 4(21) = 84

Edwin




Answer by tajar(5)

RULE to use: Translate word problem into algebraic equations.